factorize
(2x+3b)^2+(4x+6y)(a+b)+(a+b)^2
1
Expert's answer
2012-06-19T10:45:38-0400
we use formula
(s+t)^2=s^2+2*s*t+t^2
(2x+3y)^2+(4x+6y)(a+b)+(a+b)^2=(2x+3y)^2+2*(2x+3y)*(a+b)+(a+b)^2=(2x+3y+a+b)^2